2x/3 + 3y/4 = 11/12
x/3 + 7y/18 = 1/2

Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

2x/3 + 3y/4 = 11/12
x/3 + 7y/18 = 1/2

8x+9y= 12*11/12
8x+9y=11----------------1

x/3 +7y/18 =1/2
6x+7y= 1/2*18
6x+7y=9-----------------2
Multipy equation 1 by 3 and equation 2 by 4

24x+27y=33
24x+28y=36
-y=-3
y=3
Subsyitute the value of y in equation 1 an you get x=-2

Answer by stanbon(75887)

2x/3 + 3y/4 = 11/12
x/3 + 7y/18 = 1/2
-------
Multiply thru 1st equation by 12
Multiply thru 2nd equation by 18
----------------------------------------
8x + 9y = 11
6x + 7y = 9
------------------------
Multiply thru 1st equation by 6
Multiply thru 2nd equation by 8
---------------------------------------
48x + 54y = 66
48x + 56y = 72
-----------------------
Subtract 1st from 2nd1st and solve for "y":
2y = 6
y = 3
---------
Substitute into 6x+7y = 9 to solve for "x":
6x + 7*3 = 9
6x = -12
x = -2
----------------------
Cheers,
Stan H.